### v1

# A meshless method for computational electromagnetics with improved dispersion properties

Abstract
The finite difference time domain method is one of the most popular methods in computational electromagnetics. This work considers two possible ways of generalising it to a meshless setting by employing local radial basis function interpolation. The resulting methods turn out to be unstable, but can be stabilised by adding properly chosen hyperviscosity terms to the update equations. We demonstrate that the proposed meshless methods are convergent and can enjoy a decreased dispersion anisotropy compared to the finite difference time domain method.
1 Introduction
Accurate simulations of electromagnetic waves are of great interest nowadays with one of the key drivers at the moment being the field of telecommunications, as the carrier frequencies for future communication protocols (e.g. 6G) are expected to be higher, demanding physically accurate simulations that are able to describe increasingly finer geometrical objects, such as antennas. In order to have such a description, we must solve the equations that describe the physics of electromagnetism - Maxwell’s equations. Maxwell’s equations consist of a coupled system of Partial Differential Equations (PDEs) that are too complex to be solved analytically in realistic setups and a numerical procedure must be invoked.
Most of the commonly used methods in computational electromagnetics have been around for decades with several landmark papers appearing in the 1960s. In 1966, for instance, the well known Finite Difference Time Domain (FDTD) method was proposed by Yee [35]. In the same year, the Method of Moments (MoM) was popularised by Richmond [18]. The Finite Element Method (FEM), which can be traced all the way to the work of Courant in 1943 [4], has also been applied to electromagnetics by Silvester in 1969 [24]. A comprehensive review and comparison of these can be found in any of the major textbooks on computational electromagnetics, such as [5]. Among the listed methods, the FDTD is perhaps the simplest to implement and therefore a common starting point when dealing with a given problem. Its simplicity lies in the fact that the FDTD is, in essence, just the application of the usual Finite Difference Method (FDM) on a staggered grid. This approach, however, has its limitations - it is limited to regular, grid-like geometries, making it difficult to efficiently satisfy our requirement of describing fine geometrical objects. Additionally, the underlying grid makes the numerical dispersion relation anisotropic, even if the propagation medium is isotropic.
Our goal in this paper is to remedy both of these shortcomings of the FDTD by introducing a method that in a certain sense generalises it to irregular node layouts, providing a greatly increased geometrical flexibility.
For that purpose, we will adopt meshless methods, which provide the means of approximation on scattered nodes without any additional structure. Meshless or meshfree methods encompass a variety of different methods, some of which are described in, for example [9]. For our work we opt for methods based on Radial Basis Functions [3] (RBFs). RBFs have first appeared in the context of scattered data interpolation in the 1970s [13] and differential operator approximation in the 1990s [14]. In the 2000s a local RBF-based operator approximation method was introduced [30], which directly generalises the FDM, sparking an interest from the wider scientific community and resulting in a culmination of research in this field, both theoretical and experimental [31, 2].
Having the machinery of meshless methods, a generalisation of the FDTD is then straightforward - we can simply replace the Finite Differences (FD) appearing in the FDTD by their appropriate meshless variant. We will see that there are two natural ways of doing so, resulting in two different generalisations to consider.
At this point, a natural question could arise - why should one bother with a meshless variant of the FDTD, if FEM already solves Maxwell’s equations (albeit in weak form) and is known to be geometrically flexible? Our motivation for considering meshless methods over FEM is twofold. Firstly, unlike meshless methods, FEM requires an underlying triangulation, a generation of which can be tedious and cannot be fully automated, especially in higher dimensions [11]. Secondly, FEM is a fairly complex method - the attractive point of the FDTD is its simplicity and our aim is to preserve this property in the geometrically flexible version. Simplicity of the method can be of great importance when dealing with problems of increasing complexity - not only with regards to the problem geometry but also, for example, inclusion of non-homogenous or even non-linear materials. Furthermore it is also easier to build upon from an implementation perspective, in particular when it comes to parallelisation [5]. Indeed, we will see that, just like the FDTD, our proposed methods will be fully explicit.
It should be mentioned that computational electromagnetics has already been studied in a meshless context. Most of existing work has been focused on weak-form meshless methods [20, 7, 6], although some of the research has been done also using the Method of Fundamental Solution (MFS) [34], which is closely related to the previously mentioned MoM method. As mentioned, however, we are interested in strong-form methods, which resemble FDTD. The latter have also been already investigated in a meshless context [36, 22, 21] by replacing the spatial derivatives appearing in the FDTD by their meshless variants in a similar way as we intend to. However, our proposed method differs from these by two key points, which all of the listed work has in common. Firstly, the listed work staggers the $\vec{E}$ and $\vec{H}$ fields on different nodes, as in the usual FDTD. This staggering can become non-trivial in a scattered setting and would require, for example, a Voronoi tessellation, which some would argue is not truly a meshless method. The second key point is much more important - despite the methods being meshless in their formulations, they seemed to have only been tested on regular node layouts, where the methods can be much better behaved, particularly in terms of stability. In our approach we will abandon the staggering of the $\vec{E}$ and $\vec{H}$ fields and one of our goals is to demonstrate that our methods work also on scattered, irregular nodes.
The paper is structured as follows: In Section 2, the governing equations of electromagnetism are listed. In Section 3, the FDTD is briefly described, along with its previously mentioned shortcomings. In Section 4, RBF interpolation-based meshless methods are introduced. In Section 5, our proposed methods are presented and analysed. In Section 6, the dispersive properties of the proposed methods are also analysed and compared to the FDTD. We then conclude with Section 7.
2 Governing equations
In electromagnetics, the relevant quantities are the electric and magnetic fields, which we denote by $\vec{E}$ and $\vec{H}$, respectively. They are defined on some region of interest $\Omega\subset\mathbb{R}^{3}$ and are generally time-dependent:
$\vec{E},\vec{H}:\Omega\times\mathbb{R}_{\geq 0}\to\mathbb{R}^{3}.$ (1)
The two fields satisfy Maxwell’s curl equations, which in vacuum read:
$\frac{\partial\vec{H}}{\partial t}=-\frac{1}{\mu_{0}}\nabla\times\vec{E},$ (2)
$\frac{\partial\vec{E}}{\partial t}=\frac{1}{\epsilon_{0}}\nabla\times\vec{H},$
where $\epsilon_{0}$ and $\mu_{0}$ are the vacuum permittivity and permeability, respectively. Forthcoming analyses will be performed on an effectively 2-dimensional setup, where the fields $\vec{E}$ and $\vec{H}$ only depend on $(x,y)$ coordinates. For this case, it turns out [12] that the above system decouples in two systems, corresponding to the so-called TMz and TEz modes - Transverse Magnetic and Electric, respectively, in the z-direction.
TMz mode considers $\vec{E}(x,y)=(0,0,E_{z}(x,y))$, $\vec{H}=(H_{x}(x,y),H_{y}(x,y),0)$, for which the equations simplify to:
$\frac{\partial H_{x}}{\partial t}=-\frac{1}{\mu_{0}}\frac{\partial E_{z}}{\partial y},$ (3)
$\frac{\partial H_{y}}{\partial t}=\frac{1}{\mu_{0}}\frac{\partial E_{z}}{\partial x},$
$\frac{\partial E_{z}}{\partial t}=\frac{1}{\epsilon_{0}}\left(\frac{\partial H_{y}}{\partial x}-\frac{\partial H_{x}}{\partial y}\right),$
and analogously for the TEz mode with the roles of $\vec{E}$ and $\vec{H}$ interchanged. For that reason we will focus only on the TMz mode during our studies.
3 The Finite Difference Time Domain
One of the simplest and most widespread algorithms for solving the aforementioned equations is the Finite Difference Time Domain (FDTD) method, also known as the Yee algorithm. As the name implies, the algorithm discretises the curl equations using the usual Finite Difference (FD) schemes. What sets FDTD algorithm apart from the Finite Difference Method (FDM) is the use of a staggered grid (in this context known as Yee grid) - different components of $\vec{E}$ and $\vec{H}$ fields are discretised at different points in time and space. In the case of a TMz mode the field values are discretised as follows:
$E_{z}^{n,i,j}\mathrel{\mathop{:}}=E_{z}(x=i\Delta s,\ y=j\Delta s,\ t=n\Delta t),$ (4)
[…]
which is displayed on Figure 1.
Figure 1: Yee grid with a spacing of $\Delta s$. An example domain boundary $\partial\Omega$ is also displayed.
Due to the layout of the staggered grid, all of the derivatives arising in a system of equations (3) can then be discretised using central differences, resulting in an explicit scheme [^1] This manner of time discretisation is also known as interleaved leapfrog integration.:
$H_{x}^{n+1,i,j}$ $=H_{x}^{n,i,j}-S_{c}/\eta_{0}\left(E_{z}^{n,i,j+1}-E_{z}^{n,i,j}\right)$ (5)
$H_{y}^{n+1,i,j}$ $=H_{y}^{n,i,j}+S_{c}/\eta_{0}\left(E_{z}^{n,i+1,j}-E_{z}^{n,i,j}\right)$
$E_{z}^{n+1,i,j}$ $=E_{z}^{n,i,j}+\eta_{0}S_{c}\left(H_{y}^{n+1,i,j}-H_{y}^{n+1,i-1,j}-H_{x}^{n+1,i,j}+H_{x}^{n+1,i,j-1}\right),$
where $S_{c}=\frac{c_{0}\Delta s}{\Delta t}$ is the Courant number, $c_{0}^{2}=(\epsilon_{0}\mu_{0})^{-1}$ the speed of light, and $\eta_{0}^{2}=\frac{\mu_{0}}{\epsilon_{0}}$ the vacuum impedance. It can be shown that the above scheme is stable if $S_{c}\leq\frac{\sqrt{2}}{2}$. For further details on the algorithm we refer the reader to any FDTD textbook, such as [12].
3.1 Shortcomings of FDTD
The FDTD has two main shortcomings that we aim to address with the proposed methods.
The first could already be observed in Figure 1 - a grid-based method has trouble describing irregular boundaries, and the domain has to terminate in a staircase manner. Such a poor boundary description can greatly affect accuracy in more realistic geometries and often, the grid is forced to be made increasingly finer, resulting in very long computational times. Remedies for that, such as the contour FDTD exist [5], however they require manual intervention and can become cumbersome for increasingly complex and realistic domains.
The second shortcoming can be seen on Figure 2, where we simulate the effect of a Gaussian point hard [^2] Hard source at a point means we simply set the field value to a desired source function at each iteration. source, $E_{z}(n)\propto e^{-(n-100)^{2}/2}$, where $n$ is the timestep. We can see numerical dispersion appear in the form of additional, unphysical waves behind the main wave front. The interesting part is the fact that dispersion is angle dependent - with more dispersion in the $x,y$ directions and less on the $y=\pm x$ diagonals.
Figure 2: Snapshot of the FDTD simulation with the described source.
It can be derived [12] that the numerical dispersion relation in FDTD is
$\left(\frac{\Delta s}{c\Delta t}\right)^{2}\sin^{2}\left(\frac{\omega\Delta t}{2}\right)=\sin^{2}\left(\frac{\Delta sk_{x}}{2}\right)+\sin^{2}\left(\frac{\Delta sk_{y}}{2}\right),$ (6)
where the wave vector is $\vec{k}=(k_{x},k_{y})=k(\cos\phi,\sin\phi)$. As observed, the dispersion relation is anisotropic and depends on the angle of propagation $\phi$, even though the underlying physics is isotropic - the physically correct dispersion relation is well-known: $\omega=ck$. The cause of the unphysical anisotropy is the presence of a grid, which breaks the symmetry in a sense that the $x$ and $y$ directions are special, despite the fact that the curl operator appearing in Maxwell’s equations is rotationally symmetric.
It should be mentioned that an adequately dense grid can greatly reduce the effects of numerical dispersion. However, increasing the grid density also increases the computational time of the simulations and may not be viable.
4 Computation on scattered nodes
To improve the flexibility of the numerical method and ease the description of irregular boundaries, we can abandon the idea of a grid and allow the nodes to be positioned irregularly, such as in Figure 3. To obtain such a set of nodes we use a scattered node generation algorithm, described in [25], which efficiently discretises a given computational domain, producing nodes with the internodal spacing of approximately $h$ (a chosen parameter). The algorithm can also discretise more complex domains, such as CAD models [8]. We will denote the nodes of the resulting discretisation as $\mathbf{x_{i}}=(x_{i},y_{i})$, where $i$ ranges from $1$ to $N$, the discretisation size.
4.1 Interpolation on scattered nodes
In order to operate on scattered nodes, we need to invoke some tools from meshless numerical analysis. Namely, we will employ Radial Basis Functions (RBFs) to approximate our unknowns, which we now briefly describe. A more in-depth explanation can be found in [9, 16].
We start by introducing the RBF interpolation. We are looking for an approximation of a function $f:\Omega\to\mathbb{R}$ that has known values $f_{i}=f(\mathbf{x}_{i})$ on a scattered node set $X=\{\mathbf{x}_{i}\}_{i=1}^{N}$. The $N$ radial basis functions are defined by our discretisation as: $\Phi_{i}(\mathbf{x})=\varphi(\|\mathbf{x}-\mathbf{x}_{i}\|)$, for some chosen radial basic function $\varphi$. The RBF interpolant is then:
$s(\mathbf{x})=\sum_{i=1}^{N}\alpha_{i}\Phi_{i}(\mathbf{x}).$ (7)
Traditionally, $\varphi$ is chosen to be a strictly positive definite function (as are many popular choices, such as the Gaussians or the Multiquadrics). In that case, the interpolation problem can be shown to admit a unique solution [9]. In practice, to achieve a certain degree of polynomial reproduction we also add monomials to the interpolant:
$s(\mathbf{x})=\sum_{i=1}^{N}\alpha_{i}\Phi_{i}(\mathbf{x})+\sum_{j=1}^{M_{m}}\beta_{j}p_{j}(\mathbf{x}),$ (8)
[…]
$\tilde{A}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}=\begin{bmatrix}A&P\\ P^{T}&0\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}=\begin{bmatrix}f\\ 0\end{bmatrix}$, (9)
where $f=(f_{1},\dots,f_{n})^{T}$, $A_{ij}=\varphi(\|\mathbf{x}_{i}-\mathbf{x}_{j}\|)$ and $P_{ij}=p_{j}(\mathbf{x}_{i})$. Defining $\Phi(\mathbf{x})=(\Phi_{1}(\mathbf{x}),\dots,\Phi_{N}(\mathbf{x}))^{T}$ and $p(\mathbf{x})=(p_{1}(\mathbf{x}),\dots,p_{M_{m}}(\mathbf{x}))^{T}$, we can rewrite the interpolant in a more compact form:
[…]
where $\psi_{i}(x)=(\tilde{A}^{-1}(\Phi(\mathbf{x})\ p(\mathbf{x}))_{i}$, is the $i$-th cardinal function [^4] Also known as Lagrange function. for the interpolation problem and satisfies $\psi_{i}(\mathbf{x}_{j})=\delta_{ij}$.
So far, we have described a global method - the interpolation problem was solved over the whole domain. In practice, the number of nodes $N$ is big and this quickly becomes computationally inefficient. For that reason, we are usually interested in local approximations near a given point $\mathbf{x}_{c}$. In this case we simply write the interpolant (8) over a certain neighbourhood of the point $\mathbf{x}_{c}$, commonly known as its stencil. A stencil is usually taken to simply consist of $n$ ($n\ll N$) nodes that are closest to $\mathbf{x}_{c}$. An example of a stencil can be seen on Figure 3. If we denote the index set of the stencil as $S(\mathbf{x}_{c})$, the local interpolant becomes:
$s_{c}(\mathbf{x})=\sum_{i\in S(\mathbf{x}_{c})}\alpha^{c}_{i}\Phi_{i}(\mathbf{x})+\sum_{j=1}^{M_{m}}\beta^{c}_{j}p_{j}(\mathbf{x})=\sum_{i\in S(\mathbf{x}_{c})}\psi^{c}_{i}(\mathbf{x})f_{i},$ (11)
where the unknown coefficients depend on the center point $\mathbf{x}_{c}$ and can be obtained by solving a linear system, as before. It can be shown that for a sufficiently smooth $f$ the error of the interpolation scales as $\mathcal{O}(h^{m+1})$ [31].
Figure 3: An example scattered discretisation, with the nodal spacing of approximately $h$. The domain boundary $\partial\Omega$ and a stencil for a chosen node are also displayed.
4.2 Differential operator approximation
We aim to approximate a linear differential operator $\mathcal{L}$ as a linear combination of the function values in a stencil, akin to the usual finite differences:
$\mathcal{L}f(\mathbf{x}_{c})=\sum_{i\in S(\mathbf{x}_{c})}w_{i}^{\mathcal{L}}f(\mathbf{x}_{i})$ (12)
We can obtain such an approximation using a local RBF interpolant in two different ways. The usual approach is the Radial Basis Function-generated Finite Differences (RBF-FD), which simply amounts to applying the operator $\mathcal{L}$ to the interpolant:
$\mathcal{L}f(x_{c})\approx\mathcal{L}s_{c}(x_{c})=\sum_{i\in S(\mathbf{x}_{c})}\mathcal{L}(\psi_{i}^{c}(\mathbf{x}_{c}))f(\mathbf{x}_{i})=\sum_{i\in S(\mathbf{x}_{c})}w^{\mathcal{L}}_{i}f(\mathbf{x}_{i}).$ (13)
The truncation error of this approach can be shown to scale as $\mathcal{O}(h^{m+1-l})$ [31], where $l$ is the order of the highest derivatives appearing in $\mathcal{L}$.
The second approach is not as widely used in the meshless community, although it is also based on RBF interpolant and has been sucessfully applied to non-trivial problems [33, 28]. It requires us to have a way of approximating $\mathcal{L}$ with the usual finite differences - we have a finite difference stencil given by coefficients $c=(c_{1},\dots,c_{k})$ and offsets $o=(\mathbf{o}_{1},\dots,\mathbf{o}_{k})$. For example, for the central derivative in the $x$ direction we have c=(0,1,1)c=(0,-1,1) and o=((0,0),(1,0),(1,0))o=((0,0),(-1,0),(1,0)), while for the 5-point Laplace operator approximation we have c=(4,1,1,1,1)c=(-4,1,1,1,1), o=((0,0),(1,0),(1,0),(0,1),(0,1))o=((0,0),(1,0),(-1,0),(0,1),(0,-1)). Approximating the operator $\mathcal{L}$ in the finite difference manner then amounts to:
$\mathcal{L}f(\mathbf{x}_{c})\approx\delta^{-l}\sum_{i=1}^{k}c_{i}f(\mathbf{x}_{c}+\delta\mathbf{o}_{i}),$ (14)
Figure 4: A visual demonstration of the virtual stencil approach for the case of a 5-point stencil with a spacing $\delta$.
where $l$ is the order of the operator $\mathcal{L}$ and $\delta$ the spacing between the nodes. While this works well on a grid, the problem in a scattered layout is that $\mathbf{x}_{c}+\delta\mathbf{o}_{i}$ is in general not in our discretisation, so we do not have access to the value of $f$ at this point. We circumvent this by using the RBF interpolant, effectively imagining a virtual stencil around the point $x_{c}$, as seen on Figure 4:
$\mathcal{L}f(\mathbf{x}_{c})\approx\delta^{-l}\sum_{i=1}^{k}c_{i}s_{c}(\mathbf{x}_{c}+\delta\mathbf{o}_{i})=\delta^{-l}\sum_{i=1}^{k}c_{i}\sum_{j\in S(\mathbf{x}_{c})}\psi_{j}^{c}(\mathbf{x}_{c}+\delta\mathbf{o}_{i})f(\mathbf{x}_{j})=\sum_{j\in S(\mathbf{x}_{c})}w^{\mathcal{L}}_{j}f(\mathbf{x}_{j}),$ (15)
where $w^{\mathcal{L}}_{j}=\delta^{-l}\sum_{i=1}^{k}c_{i}\psi_{j}^{c}(\mathbf{x}_{c}+\delta\mathbf{o}_{i})$, are the differentiation weights, expressing the approximation of the differential operator $\mathcal{L}$ and can again be understood as the generalised finite difference coefficients, with the offsets given by the stencil nodes. The virtual stencil spacing $\delta$ is a parameter of the method an usually expressed in the units of the internodal spacing $h$ as $\delta=\sigma h$. Experiments show that it is desirable to have $\sigma<1$ [19, 33]. As far as the authors know, the second approach does not have a well-established name and we will refer to it as Radial Basis Function-generated Virtual Finite Differences (RBF-VFD) in this paper. The order of the RBF-VFD method can be easily deduced: The interpolation step has order $\mathcal{O}(h^{m+1})$ and the interpolant is divided by $h^{l}$, bringing the order to $\mathcal{O}(h^{m+1-l})$. The second step, the application of the chosen finite difference stencil has a fixed, known order $\mathbf{ord}$. The total order is then simply $\mathrm{min}(m+1-l,\mathbf{ord})$.
4.3 Details of our setup
In our work, we have exclusively used Polyharmonic Splines (PHS) as our chosen RBF. Concretely, we use radial cubics $\varphi(r)=r^{3}$, which are conditionally positive definite of order $2$ and thus require the addition of monomials up to at least $m=1$ [9]. It has been shown that such a combination can maintain the approximation power of monomials, while offering increased stability if the stencil size is chosen to be at least $n=2M_{m}$ [2], which will be our default choice as well.
For meshless numerical analysis, we use an open source C++ library Medusa [26]. All of our source codes and data are publicly available on a git repository [^5] https://gitlab.com/e62Lab/2025_p_meshless_computational_electromagnetics.
5 Generalising the FDTD to a meshless setting
Having the tools of numerical analysis on scattered nodes behind us, we can now proceed to implementing a variant of the FDTD in a meshless setting.
Firstly, the previously mentioned shortcomings of FDTD arise due to the grid discretisation in space. Time direction has a trivial geometry and poses no issue, so we will keep the staggering of fields in time and the same time evolution method - the interleaved leapfrog.
The problematic part is the spatial discretisation, which will be given by scattered nodes. We have decided to discard the idea of staggering the $\vec{H}$ and $\vec{E}$ fields at different points, as this could become nontrivial in complex settings (even more so in higher dimensions). Instead, we simply take the FDTD update equations (5) and replace the spatial derivatives (central differences) by their suitable meshless substitutes:
$H_{x}(x_{i},y_{i},n+1)=H_{x}(x_{i},y_{i},n)-\frac{\Delta t}{\mu_{0}}D^{\partial_{y}}(E_{z}(x_{i},y_{i},n)),$ (16)
$H_{y}(x_{i},y_{i},n+1)=H_{y}(x_{i},y_{i},n)+\frac{\Delta t}{\mu_{0}}D^{\partial_{x}}(E_{z}(x_{i},y_{i},n)),$
$E_{z}(x_{i},y_{i},n+1)=E_{z}(x_{i},y_{i},n)+\frac{\Delta t}{\epsilon_{0}}\left(D^{\partial_{x}}H_{y}(x_{i},y_{i},n+1)-D^{\partial_{y}}H_{x}(x_{i},y_{i},n+1)\right),$
where we have introduced the notation $D^{\mathcal{L}}f(\mathbf{x}_{i})=\sum_{j\in S(\mathbf{x}_{i})}w_{j}^{\mathcal{L}}f_{(}\mathbf{x}_{j})$ and the two approaches that we will consider differ in how the differentiation weights $w_{j}^{\mathcal{L}}$ are calculated.
In the first approach, we will calculate the differentiation weights using RBF-VFD, where the finite difference stencils in question are simply the central differences. This approach seems to be as close to the Yee grid as we can get to in the meshless setting. We have opted for $\delta=0.5h$ as our virtual stencil spacing of choice, which is within the previously mentioned guidelines and mirrors the spacing of the original Yee grid, where the $\vec{E}$ and $\vec{H}$ are also offset by the same amount.
In the second approach we will apply RBF-FD to obtain the differentiation weights, which directly generalises the usual FDTD [15].
In order to match the order of FDTD, we calculate the derivative weights using $m=2$.
The above scheme remains explicit and thus efficient, as the differentiation weights are only computed once at the beginning of the simulation and can be reused. We gain additional flexibility of the method due to the meshless nature of the differential operators.
However, our attempt at a more geometrically flexible version of the FDTD comes at a cost in terms of the method stability.
In FDTD it can be shown that the method is stable if the Courant number satisfies $S_{c}:=\frac{c_{0}\Delta t}{\Delta s}\leq S_{c}^{\mathrm{max}}$, where $S_{c}^{\mathrm{max}}$ equals $\frac{1}{\sqrt{2}}$ for the two-dimensional case with similar stability conditions on the Courant number available also for other regular node layouts [12]. However, in a meshless setting, the nodes are scattered and such an estimate is not easily reachable. We have attempted to stabilise the method by computing at an increasingly lower $S_{c}$ with no success [^6] For the forthcoming analyses, we have set $S_{c}=(10\sqrt{2})^{-1}$., hinting at an inherent instability of the method. Refer to the Appendix for an illustration of why instability may be expected. To remedy these problems and make our proposed methods useful, a stabilisation procedure is needed.
5.1 Stabilising the method
Instabilities in explicit numerical schemes are a well-known problem and subject to a plethora of research in the scientific community. Many different approaches to stabilisation have been proposed over the years, such as penalty schemes or upwinding, to name a few.
We have opted for an approach that has been gaining popularity in the meshless community lately: Hyperviscosity (HV) [10]. It can be effective and is in principle very easy to implement, as it amounts to simply adding a higher order derivative term to our system:
$\frac{\partial H_{x}}{\partial t}=-\frac{1}{\mu_{0}}\frac{\partial E_{z}}{\partial y}+\gamma_{1}\Delta^{\alpha_{1}}H_{x},$ (17)
$\frac{\partial H_{y}}{\partial t}=\frac{1}{\mu_{0}}\frac{\partial E_{z}}{\partial x}+\gamma_{2}\Delta^{\alpha_{2}}H_{y},$
$\frac{\partial E_{z}}{\partial t}=\frac{1}{\epsilon_{0}}\left(\frac{\partial H_{y}}{\partial x}-\frac{\partial H_{x}}{\partial y}\right)+\gamma_{3}\Delta^{\alpha_{3}}E_{z},$
where $\Delta$ is the Laplace operator. For further intuition on how HV stabilises a given method, an interested reader may refer to [23, 10].
What remains is a determination of parameters $\gamma_{i},\alpha_{i}$ that should be set such that the HV term stabilises the solution, while not affecting the physically relevant solution components too much. We first perform some simplifications that turned out to work well. We set:
$\alpha_{1}=\alpha_{2}=\alpha_{3}=\alpha,$ (18)
$\gamma_{1}=\gamma_{2}=\frac{\gamma}{\mu_{0}},$
$\gamma_{3}=\frac{\gamma}{\epsilon_{0}}.$
These choices are motivated by the fact that we would like to have as few parameters as possible, while still taking into account the asymmetry between $\vec{H}$ and $\vec{E}$.
Additionally, we follow the hyperviscosity guidelines described in [23] that set $\gamma=c\ (-1)^{\alpha-1}h^{2\alpha}$, for some constant $c$. We will discretise the newly introduced hyperviscosity operator $\gamma_{i}\Delta^{\alpha}$ using RBF-FD, resulting in modified update equations of the stabilised method:
$H_{x}(x_{i},y_{i},n+1)=H_{x}(x_{i},y_{i},n)-\frac{\Delta t}{\mu_{0}}\left(D^{\partial_{y}}(E_{z}(x_{i},y_{i},n))-\gamma D^{\Delta^{\alpha}}H_{x}(x_{i},y_{i},n)\right),$
$H_{y}(x_{i},y_{i},n+1)=H_{y}(x_{i},y_{i},n)+\frac{\Delta t}{\mu_{0}}\left(D^{\partial_{x}}(E_{z}(x_{i},y_{i},n))+\gamma D^{\Delta^{\alpha}}H_{y}(x_{i},y_{i},n)\right),$ (19)
$E_{z}(x_{i},y_{i},n+1)=E_{z}(x_{i},y_{i},n)+\frac{\Delta t}{\epsilon_{0}}\left(D^{\partial_{x}}H_{y}(x_{i},y_{i},n+1)-D^{\partial_{y}}H_{x}(x_{i},y_{i},n+1)+\gamma D^{\Delta^{\alpha}}E_{z}(x_{i},y_{i},n)\right).$
Note that, for simplicity, we will use RBF-FD for the HV part even if the $\partial_{x},\partial_{y}$ operators are discretised using RBF-VFD. The rationale is that the HV term is merely a stabilisation procedure and should not be responsible for other properties of the method. Following also the guidelines presented in [32], we choose $\varphi(r)=r^{2\alpha+1}$, $m=2\alpha$ and $n=2M_{m}+1$ for the RBF-FD discretisation of the HV operator.
The values of the parameters $\alpha$ and $c$ are determined by a parameter sweep - we solve a given problem with some chosen parameters and observe its stability properties.
To quantify (in)stability of a method, we track the behaviour of energy over time for different values of $c$ and $\alpha$. The formula of energy of the electromagnetic field is known:
$U(t)$ $\propto\int_{\Omega}\|\vec{E}(t)\|^{2}+\eta_{0}^{2}\|\vec{H}(t)\|^{2}\textup{d}\Omega$
$\approx h^{2}\sum_{i=1}^{N}\left(E_{z}(x_{i},y_{i},t)^{2}+\eta_{0}^{2}(H_{x}(x_{i},y_{i},t)^{2}+H_{y}(x_{i},y_{i},t)^{2})\right),$
where the integral was approximated with a sum similarly as in [31]. Due to staggering in time, we do not have access to the $\vec{E}$ and $\vec{H}$ fields at the same time points. In order to calculate the energy at an integer multiple of $\Delta t$, we must average the $\vec{H}$ fields (which are otherwise only defined at odd multiples of $\Delta t/2$). We should emphasize that an analytical solution of the problem is not needed to compute $U(t)$. In practice, we can observe the relative energy $u(t)=U(t)/U(0)$, which should remain approximately $1$ for a conservative problem setup.
As an example, we solve Maxwell’s equations on a domain $\Omega=[0,200]^{2}$, with periodic boundary conditions (emulating an unbounded domain) and initial condition given by:
$E_{z}(x,y,t=0)=e^{-\frac{(x-100)^{2}}{2\cdot\sigma^{2}}},$ (20)
$0,\displaystyle H_{x}(x,y,t=-5\Delta t)=$0,
[…]
which is a Gaussian pulse, that should analytically move at a speed $c_{0}$ in the $x$ direction without changing its shape. We choose $\sigma=30$ as its width.
Figure 5: The behaviour of relative energy $u(t)$ over time for the considered test case for different hyperviscosity parameters $\alpha$ and $c$.
The time dependence of $u(t)$ for different stabilisation parameters is displayed in Figure 5. We can observe that $c$ has a major effect on the stability of the method and that a well-chosen value of $c$ and $\alpha$ leads to an approximate conservation of energy throughout the simulation. A more complete picture of the effect of $c$ on stability can be seen in Figure 6, where the energy after $10000$ timesteps with respect to $c$ is shown. Here, a common pattern in HV can be observed: There exists an interval of suitable values of $c$, for which the solution is stable. For $c$ outside this interval, the solution diverges. Note that stable schemes may also be dissipative, which in our case happens at $\alpha=1$ and is undesirable. In more complex settings the interval of $c$, for which the method is stable, can be very narrow and dissipation commonly arises, even in $\alpha>1$ case. For that reason, complex settings may require a very fine parameter sweep and for practical reasons, we may to resort to the bisection approach to efficiently calculate a suitable value of $c$.
For the remainder of this section, we have settled on $\alpha=2$ and $c=10^{-4}$, as it is a stable and non-dissipative combination. Generally, we prefer a lower $\alpha$ if possible, as the stencil size and therefore the computational complexity of the method grows with $\alpha$.
Figure 6: Relative energy after $10000$ timesteps with respect to $c$ for different values of the parameter $\alpha$. The value $\infty$ represents the divergence of the solution.
5.2 Convergence of the solution
To end this section, let us verify that the proposed scheme is convergent. In the above setup $H_{x},H_{y}$ were chosen such that the wave propagates into the positive $x$ direction (the propagation direction is given by $\vec{E}\times\vec{H}$). Since the problem is independent of $y$, it is effectively 1-dimensional and as $E_{z}$ satisfies the wave equation, the solution is given by the d’Alembert formula and equals:
$E^{a}_{z}(x,y,t)=E_{z}(x-c_{0}t,y,0)$ (21)
$H^{a}_{x}(x,y,t)=0,$
$H^{a}_{y}(x,y,t)=-E_{z}(x-c_{0}t,y,0)/\eta_{0},$
where we have added the superscript $a$ to denote the analytical solution and differentiate from the numerically obtained solution. This allows us to verify that our stabilised method is indeed convergent. We check this by reducing the spatial and time discretisation, while keeping the Courant number $S_{c}=\frac{1}{10\sqrt{2}}$ constant. To study the convergence we will define the following two $\ell^{\infty}$ errors:
$\|e\|^{E}_{\infty}\mathrel{\mathop{:}}=\mathrm{max}_{i=1}^{N}\left(\|\vec{E}(\mathbf{x}_{i})-\vec{E}^{a}(\mathbf{x}_{i})\|_{\infty}\right),$ (22)
$\|e\|^{H}_{\infty}\mathrel{\mathop{:}}=\mathrm{max}_{i=1}^{N}\left(\|\vec{H}(\mathbf{x}_{i})-\vec{H}^{a}(\mathbf{x}_{i})\|_{\infty}\right).$
The results are displayed in Figure 7, where we can see that both methods are convergent with similar errors behaviour for both methods as $h\to 0$. This dependence maintains its shape also if we consider slightly different discretisations by modifying the random seed in our node generation algorithm [25]. Ideally, a convergence rate of $\propto h$ is expected to occur, however, a slightly slower convergence is not surprising due to hyperviscosity.
Figure 7: $\|e\|^{H}_{\infty}$ and $\|e\|^{E}_{\infty}$ errors with respect to the discretisation distance $h$.
To summarise, we now have a stable method that can operate on scattered nodes. With this we have addressed the first shortcoming of the FDTD and we now turn to the second - anisotropical behaviour of the dispersion relation.
6 Dispersion
Having developed a working meshless algorithm, we now turn to its dispersion properties.
To evaluate the magnitude of dispersion of our methods, we consider a similar test case as in Section 5.1, except that we use $\sigma=2$ - a very narrow Gaussian that is difficult to resolve at $h=1$ and will push the method to its limits inducing visible effects of dispersion. As the setup is now different, the previously calculated HV parameters are not suitable and we recompute them in the same manner as before. As the newly considered setup is much more unstable, a naive parameter sweep is slow and we have computed the parameters using the bisection approach. We have settled for $\alpha=2$ and $c\approx 10^{-4.76}$ for the RBF-FD and $c\approx 10^{-5.02}$ for RBF-VFD, which maintained $u(t)\in[0.99,1.01]$ even after $10000$ timesteps.
To analyse dispersion we will turn to Fourier analysis, which works well on uniformly spaced nodes. For that reason, we perform an extra step when discretising $0,200]^{2} -$ first, we discretise the line $y=100$ with equispaced nodes $\{\mathbf{x}^{e}_{i}\}_{i}$, at a distance of $h=1$ apart. These are then used as seed nodes in the same discretisation algorithm to obtain the complete discretisation of $0,200]^{2}$, effectively resulting in a scattered node layout, where some points are guaranteed to be equispaced on a line.
Figure 8 shows the snapshot of the simulation after $2820$ timesteps (approximately equal to one passage of the wavefront across the whole domain) for different method parameters, where several observations can be made. First, we can see that increasing the stencil size [^7] Note that this refers to the stencil size of the original, unstabilised parts of the update equations. For stabilisation, we continue to adhere to the previous guidelines. improves the dispersion properties of the solution. In fact, for this particular case, the smallest recommended stencil size $n=13$ is very sensitive to the properties of the discretisation, resulting in numerical artifacts in the solution. We do not yet fully understand the reasoning behind this behaviour, although some ideas on how to analyse it will be mentioned in the conclusion. A second thing to observe is that, perhaps surprisingly, dispersion seems much more prominent in the RBF-VFD approach compared to the RBF-FD.
Figure 8: Snapshot of the solution for different stencil sizes and methods.
To put our observations in a formal setting, let us analyse the spectral picture of the solution. Running the simulation as before, we sample the solution values $E_{z}(\mathbf{x}^{e}_{i},n\Delta t)$ and perform a two-dimensional Fourier transform to obtain the corresponding values in frequency space $E_{z}(k,\omega)$, where the allowed discrete values for $k$ are $k_{i}=\frac{2\pi i}{N_{x}}$, for $0\leq i<N_{x}$, with $N_{x}=200$ (number of points on the discretised line) and similarly $\omega_{j}=\frac{2\pi j}{15N_{t}}$ with $0\leq j<N_{t}$ with $N_{t}=188$ (running the simulation for $2820$ timesteps and saving every 15th snapshot). $k$ are given in units of $h^{-1}]$, while the units of $\omega$ are $(\Delta t)^{-1}]$.
Before transforming, Hann window was applied in the time direction to prevent spectral leakage due to non-periodicity (in time) of the solutions. The results can be seen in Figure 9, where the absolute value of the individual spectral components $E_{z}(k_{i},\omega_{j})$ is displayed. Physically, we would expect all the modes to propagate at the same speed, that is, that the magnitude of $E_{z}(k,\omega)$ is large only when $\omega=c_{0}k$, where $c_{0}=\frac{1}{10\sqrt{2}}$ in our units. A black, dashed line representing this physical dispersion relation is also displayed in the same Figure. We can clearly see the different behaviour of the cases with the smallest stencil size ($n=13$), where the linear dispersion relation breaks down already at a relatively low frequency. This is greatly improved as stencil size is increased (note the logarithmic scale - the off-diagonal components are multiple orders of magnitude smaller). As already hinted by our previous simulation, we can observe that the RBF-VFD approach further deviates from the expected dispersion relation compared to the RBF-FD.
Figure 9: $\ln|E_{z}(k,\omega)|$ for the considered cases. The $x$ and $y$ axes correspond to wavenumber and frequency, as indicated in the text. The expected dispersion relation is denoted by the black, dashed line.
For the purposes of visualisation, we would like to summarise the amount of error in the numerical dispersion with a single quantity. We first define a numerical propagation velocity $c(\omega_{j})$ by selecting for each discrete frequency $\omega_{j}$ the corresponding wave vector $k_{i}$ that is most represented:
$i(j)$ $\mathrel{\mathop{:}}=\mathrm{argmax}_{0\leq i<N_{x}}|E(k_{i},\omega_{j})|$ (23)
$c(\omega_{j})$ $\mathrel{\mathop{:}}=\frac{\omega_{j}}{k_{i(j)}}$
The absolute error in the propagation velocity is then
$\mathbf{err}(\omega_{j})\mathrel{\mathop{:}}=\left|c(\omega_{j})-\frac{1}{10\sqrt{2}}\right|$ (24)
As our quantity, summarising the amount of deviation from the numerical dispersion relation, we simply take the sligthly modified $\ell^{1}$ norm of $\mathbf{err}$:
$\|\mathbf{err}\|_{1}\mathrel{\mathop{:}}=\sum_{j=0}^{45}|\mathbf{err}(\omega_{j})|,$ (25)
where we have only summed up to $j=45$, as higher values are not visibly present in the spectrum and would distort the results. This quantity is displayed in Figure 10 with respect to the stencil size and matches our previous observations - the error is large at small stencil sizes and is reduced as the stencil size is increased. From this point onward, we will work with $n=60$ as our stencil size of choice for both methods.
Figure 10: Stencil size dependence of the chosen dispersion error.
Having delved into the dispersive properties of our presented methods it is time to introduce angular dependence and compare it to the FDTD. For this, we modify our setup yet again - we again start with $0,200]^{2}$ and a line $y=100$, however, we now rotate it by a chosen angle $\phi$, and afterwards complete the discretisation, again using the (now rotated) line as seed nodes. The initial condition is the same pulse wave as before, moving along the square (in the $\phi$ direction):
[…]
Such a setup allows us to maintain the periodic boundary condition and perform the same Fourier analysis as before, the difference being that the pulse now propagates at an angle $\phi$ with respect to the $xy$ coordinate system.
The results of the simulation for two extreme angles $\phi=0$ and $\phi=\frac{\pi}{4}$ are seen in Figure 11. Note that, while for the $\phi=0$ case, we can implement the FDTD on $0,200]^{2}$ as before, the rotated case was implemented on $0,144\sqrt{2}]^{2}$, rotated by $\frac{\pi}{4}$, as we cannot make the grid periodic otherwise. Qualitative results should be independent of the slight change in the domain size, as the width of the pulse remains the same. We can observe that the RBF-FD case is not affected by the orientation, which is what we initially hoped for, as the stencil has no directional bias. This directional independence was not a-priori obvious, however: In the RBF-FD approach we still compute the derivatives in the $x$ and $y$ direction separately and do not manually account for the fact that the curl operator is rotationally invariant.
On the contrary, the dispersion properties of the FDTD are expectedly highly angle-dependent. Interestingly, the same holds for RBF-VFD, implying that having a stencil with no directional bias is not sufficient for isotropy and the anisotropy can arise from the manner in which differentiation weights are computed.
Figure 11: Snapshots of the simulation after one passage through the domain for the two extremal angles.
We conclude the section by plotting $\|\mathrm{err}\|_{1}(\phi)$ in Figure 12, where the RBF-FD and RBF-VFD curves were computed in the same manner as before, while for the FDTD we calculated the error from Equation (6). The latter plot again shows that the RBF-FD approach exhibits a much lower degree of dispersion anisotropy, compared to the RBF-VFD and the FDTD approaches, which behave similarly.
Figure 12: Dispersion error dependence on the propagation angle $\phi$.
7 Conclusion
The starting point of our research was a seemingly simple question - How can the FDTD method be generalised to a scattered setting and would this help with the dispersion anisotropy? The first question was answered affirmatively with the use of RBF-based meshless methods - we have performed a straightforward generalisation of the FDTD by simply replacing the derivative operators appropriately. We have considered two approaches, which a-priori both seemed equally promising: The usual RBF-FD and the virtual stencil approach, in this paper referred to as RBF-VFD. Unlike FDTD, which enjoys a simple stability criterion, major issues were found regarding stability of the proposed methods and we have demonstrated that these issues can be overcome by employing hyperviscosity stabilisation, describing a simple framework on how the appropriate HV parameters should be selected.
We then analysed the dispersion properties and found that RBF-FD possesses a parameter that can control the amount of dispersion - the stencil size. Moving to the angular dependence of the dispersion, we have answered affirmatively also to the second part of our question - meshless methods can eliminate dispersion anisotropy, which in our example happened with the RBF-FD method. However, this is not always the case, as demonstrated by the RBF-VFD approach.
Ultimately, we have seen that RBF-FD performs better than FDTD in this sense and can be considered as a promising candidate for a meshless generalisation of the FDTD.
Several important points were glossed over in our research that should be addressed if the proposed methods are ever to be used in an industrial setting. First, our analysis was limited to an effectively one-dimensional case. As part of our current research, we are applying those same methods to tackle different cases, such as scattering on an irregular body or some 3D problems. This also includes analyses of the behaviour of the method, when applied to cases with spatially varying discretisation density.
An important advantage of the FDTD method that was not given much attention is its intrinsic divergence conservation [12]. In this aspect, the RBF-VFD, being in some sense closer to the FDTD, could perform better compared to the RBF-FD approach, especially if paired with the so-called divergence-free RBFs [17], paving the way for potential future work in that direction.
Finally, our observations of the stencil dependence of dispersion remain unexplained. However, we strongly suspect that the reason behind it can be connected to the recent developments in the RBF-FD method [31], where it was shown that the basis of our interpolant in the global sense is discontinuous, with the jumps decreasing with the stencil size. Further work on this matter could put our observations in a concrete, mathematical setting.
Acknowledgements
The authors would like to acknowledge the financial support from the Slovenian Research and Innovation Agency (ARIS) in the framework of the research core funding No. P2-0095, The Young Researcher program PR-12347 and research project No. J2-3048.
Furthermore, the authors would like to thank the following individuals, listed in alphabetical order, for their assistance with various aspect of this study: Vito Levstik, Vagif Osmanov, Adrijan Rogan, Miha Rot and Žiga Vaupotič.
8 Appendix: Naive instability on scattered nodes
Here we aim to provide some intuition on why a naive approach to solving Maxwell’s equations on scattered nodes without any stabilisation fails.
For convenience, let us repeat the update equations from Section 5:
$H_{x}(x_{i},y_{i},n+1)=H_{x}(x_{i},y_{i},n)-\frac{\Delta t}{\mu_{0}}D^{\partial_{y}}(E_{z}(x_{i},y_{i},n)),$ (27)
$H_{y}(x_{i},y_{i},n+1)=H_{y}(x_{i},y_{i},n)+\frac{\Delta t}{\mu_{0}}D^{\partial_{x}}(E_{z}(x_{i},y_{i},n)),$
$E_{z}(x_{i},y_{i},n+1)=E_{z}(x_{i},y_{i},n)+\frac{\Delta t}{\epsilon_{0}}\left(D^{\partial_{x}}H_{y}(x_{i},y_{i},n+1)-D^{\partial_{y}}H_{x}(x_{i},y_{i},n+1)\right),$
where $D^{\mathcal{L}}f(x,y)=\sum_{j}w^{\mathcal{L}}_{j}f(x_{j},y_{j})$ is an approximation of a differential operator, given as a linear combination of function values in a stencil.
A straightforward way to analyse the stability of a method in FD schemes is by the means of Von Neumann stability analysis [27]. Note that Von Neumann analysis assumes translational invariance and hence applies to the usual grid-based schemes. In a meshless setting it should be used with extreme caution and not as a robust tool to assess the stability, perhaps providing only a very localised picture of modes that are amplified after a single iteration. Nevertheless, a variant of Von Neumann analysis has been used by different authors in the meshless community with moderate success [23, 29].
By Von Neumann we take an ansatz888a general Fourier component has $i(k_{x}x+k_{y}y)$ in the exponential. However, we will see that unstable modes arise already in this special case.:
$H_{x}^{n}(x,y)$ $=H_{x0}^{n}e^{ik(x+y)}$ (28)
$H_{y}^{n}(x,y)$ $=H_{y0}^{n}e^{ik(x+y)}$
$E_{z}^{n}(x,y)$ $=E_{z0}^{n}e^{ik(x+y)}$
Consdering the update equations (27) for a single point $(x_{i},y_{i})$, plugging in the ansatz and dividing by $e^{ik(x+y)}$ we get:
$H_{x0}^{n+1}=H_{x0}^{n}-c_{1}w_{y}E_{z0}^{n}$ (29)
$H_{y0}^{n+1}=H_{y0}^{n}+c_{1}w_{x}E_{z0}^{n}$
$E_{z0}^{n+1}=E_{z0}^{n}+c_{2}\left(H_{y0}^{n+1}w_{x}-H_{x0}^{n+1}w_{y}\right)$
where we have introduced the notation:
$c_{1}\mathrel{\mathop{:}}=\frac{\Delta t}{\mu_{0}}$ (30)
$c_{2}\mathrel{\mathop{:}}=\frac{\Delta t}{\epsilon_{0}}$
$w_{x}\mathrel{\mathop{:}}=\sum_{j}w_{j}^{\partial_{x}}e^{ik(x_{j}-x+y_{j}-y)}$
$w_{y}\mathrel{\mathop{:}}=\sum_{j}w_{j}^{\partial_{y}}e^{ik(x_{j}-x+y_{j}-y)}$
Rewriting in matrix form, we get:
$\begin{pmatrix}1&0&0\\ 0&1&0\\ w_{y}c_{2}&-w_{x}c_{2}&1\end{pmatrix}\begin{pmatrix}H_{x0}^{n+1}\\ H_{y0}^{n+1}\\ E_{z0}^{n+1}\end{pmatrix}=\begin{pmatrix}1&0&-c_{1}w_{y}\\ 0&1&c_{1}w_{x}\\ 0&0&1\end{pmatrix}\begin{pmatrix}H_{x0}^{n}\\ H_{y0}^{n}\\ E_{z0}^{n}\end{pmatrix}$ (31)
or, equivalently:
$\begin{pmatrix}H_{x0}^{n+1}\\ H_{y0}^{n+1}\\ E_{z0}^{n+1}\end{pmatrix}=\begin{pmatrix}1&0&-c_{1}w_{y}\\ 0&1&c_{1}w_{x}\\ -c_{2}w_{y}&c_{2}w_{x}&1+c_{1}c_{2}(w_{x}^{2}+w_{y}^{2})\end{pmatrix}\begin{pmatrix}H_{x0}^{n}\\ H_{y0}^{n}\\ E_{z0}^{n}\end{pmatrix}$ (32)
For stability, it is necessary that the spectral radius of the matrix in the last equation is below unity. We can calculate the eigenvalues to equal:
$\lambda_{1}$ $=1$ (33)
[…]
We can expect unstability if one of the $\lambda_{2,3}$ has modulus greater than unity. Denoting $\xi\mathrel{\mathop{:}}=c_{1}c_{2}(w_{x}^{2}+w_{y}^{2})$ we can write:
$\lambda_{2,3}=1+0.5\left(\xi\pm\sqrt{\xi^{2}+4\xi}\right)$ (34)
Observe that the product of the two equals unity:
$\lambda_{2}*\lambda_{3}=(1+0.5\xi)^{2}-0.5^{2}(\xi^{2}+4\xi)=1$ (35)
Therefore, $\lambda_{2}=1/\lambda_{3}$. From this it follows that if $|\lambda_{2}|<1$ then $|\lambda_{3}|>1$ and an unstable mode has been found. We must only exclude the possiblity of $|\lambda_{2}|=|\lambda_{3}|=1$. However, in this case we can write $\lambda_{2}=e^{i\alpha},\lambda_{3}=e^{-i\alpha}$ and calculate:
$2+\xi=\lambda_{2}+\lambda_{3}=e^{i\alpha}+e^{-i\alpha}=2\cos(\alpha).$ (36)
We notice that the case $|\lambda_{2}|=|\lambda_{3}|=1$ is only possible if $\xi$ is real, which is something that is very unlikely to happen in an irregular stencil (real or imaginary parts in $w_{x}$$w_{y}$ would have to exactly cancel out).
As emphasised earlier, this is by no means a formal proof of the instability of any meshless approach, but instead merely an intuition as to where the locally unstable modes may come from. The analysis was limited to an interleaved leapfrog time-stepping, with no field staggering and a general expression of a derivative in a stencil. This includes our RBF-FD and RBF-VFD methods.